OpenStudy (anonymous):

Bradley and Kelly are out flying kites at a park one afternoon. A model of Bradley and Kelly’s kites are shown below on the coordinate plane as kites BRAD and KELY respectively.

OpenStudy (anonymous):

They are similar because Segment BR to segment DB is 1:3 and Segment KE to segment YK is 1:3 They are similar because Segment BR to segment DB is 1:2 and Segment KE to segment YK 1:2. They are not similar because Segment BR to segment DB is 1:2 and Segment KE to segment YK is 1:3. They are not similar because Segment BR to segment DB is 1:3 and Segment KE to segment YK is 1:2.

OpenStudy (anonymous):

hey:) can you help me ?

OpenStudy (anonymous):

i think the answer is c

OpenStudy (goformit100):

YEAH

jimthompson5910 (jim_thompson5910):

First you need to find the lengths of BR and DB

OpenStudy (anonymous):

5,6

OpenStudy (anonymous):

and 1,4

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

2nd option is the answer

jimthompson5910 (jim_thompson5910):

do you know the distance formula?

OpenStudy (anonymous):

They are similar because Segment BR to segment DB is 1:2 and Segment KE to segment YK 1:2.

OpenStudy (anonymous):

BR/DB=KE/YK=1/2

OpenStudy (anonymous):

oh ok thank you i have another one

jimthompson5910 (jim_thompson5910):

no, the answer is actually They are not similar because Segment BR to segment DB is 1:2 and Segment KE to segment YK is 1:3. so C is the answer you basically find the lengths of BR, DB, KE, and YK. Then use compute the ratios above and you'll see that they aren't equal

jimthompson5910 (jim_thompson5910):

use them to compute*

OpenStudy (anonymous):

didn't read the options properly thought...every option mentioned about similarity so calculated BR/DB=\[\sqrt{5}/2\sqrt{5}=1/2\]...so assumed that for similarity KE/YK=1/2....sorry should hv read it more carefully

OpenStudy (anonymous):

oh ok thank you so much :)